I combust 22 g of propane (C3H8) in the presence of 96 g of oxygen in a strong, sealed, 100 liter ridgid container. What is the total pressure in the container if, after combustion, the temperature is 400 K?
1
Expert's answer
2011-03-04T14:09:52-0500
C3H8 + 5O2 = 3CO2 + 4H2O
At 400 K the water is in gaseous state.
M(C3H8) = 3*12 + 8 = 44 g/mol
M(O2) = 2*16 g/mol = 32 g/mol

ν(C3H8) = 22/44 =0.5 mol;
ν(O2) = 22/44 =0.5*5 = 2.5 mol.
ν(CO2) = 0.5 * 3 = 1.5 mol
ν(H2O) = 0.5 * 4 = 2 mol.
there is ν(O2) = 3 - 2.5 = 0.5 mol is left after combustion.

According to the law

pV = νRT

the pressure would be p = νRT/V = (0.5 +2 + 1.5)* 8.31* 400K/0.1 = 132,960 Pa.